father's age is three times the sum of ages of his two childen. after 5 year his will be twice the sum of ages of the son. find the present ages of father and son.
Answer by ankor@dixie-net.com(22740)

father's age is three times the sum of ages of his two children.
after 5 years his will be twice the sum of the ages of the son.
find the present ages of father and son.
:
Let a = age of the son
Let b = age of the other child
Let f = fathers age
:
"father's age is three times the sum of ages of his two children."
f = 3(a+b))
f = 3a + 3b
:
"after 5 years his will be twice the sum of ages of the son. "
f + 5 = 2(a + 5)
f + 5 = 2a + 10
f = 2a + 10 - 5
f = 2a + 5
replace f with (3a+3b)
3a + 3b = 2a + 5
3a - 2a = - 3b + 5
a = -3b + 5
Only one positive integer solution; b = 1, (other child's age)
a = -3(1) + 5
a = 2 yrs is the son's present age
then
f = 3(2+1)
f = 9 yrs is father's age, doesn't make sense, but if fits
:
check in the statement
"after 5 year his will be twice the sum of ages of the son."
9 + 5 = 2(2+5)
14 = 2(7)